# Think Twice before Adaptation: Improving Adaptability of DeepFake Detection via Online Test-Time Adaptation

### v1

@@ Abstract
Deepfake (DF) detectors face significant challenges when deployed in real-world environments, particularly when encountering test samples deviated from training data through either postprocessing manipulations or distribution shifts. We demonstrate postprocessing techniques can completely obscure generation artifacts presented in DF samples, leading to performance degradation of DF detectors. To address these challenges, we propose Think Twice before Adaptation (T2A), a novel online test-time adaptation method that enhances the adaptability of detectors during inference without requiring access to source training data or labels. Our key idea is to enable the model to explore alternative options through an Uncertainty-aware Negative Learning objective rather than solely relying on its initial predictions as commonly seen in entropy minimization (EM)-based approaches. We also introduce an Uncertain Sample Prioritization strategy and Gradients Masking technique to improve the adaptation by focusing on important samples and model parameters. Our theoretical analysis demonstrates that the proposed negative learning objective exhibits complementary behavior to EM, facilitating better adaptation capability. Empirically, our method achieves state-of-the-art results compared to existing test-time adaptation (TTA) approaches and significantly enhances the resilience and generalization of DF detectors during inference. Code is available here.
Keywords deepfake detection, test-time adaptation, generalization, adaptability
@@ 1 Introduction
Recently, Generative Artificial Intelligence (GenAI) has been used to generate DFs for malicious purposes, such as impersonation [^1] Finance worker pays out \$currency-dollar\\$\$25 million after video call with deepfake ‘chief financial officer’ and disinformation spread [^2] AI-faked images of Donald Trump’s imagined arrest swirl on Twitter, raising concerns about privacy and security. Several DF detection approaches have been proposed to mitigate these negative impacts [1]. Despite advances, deploying these systems in real-world environments presents two critical challenges. First, in practice, adversaries can strategically apply previously unknown postprocessing techniques to DF samples at inference time, completely obscuring the generation artifacts [2] and successfully bypassing detection systems. Second, real-world applications are frequently exposed to test samples drawn from distributions that deviate substantially from the training data distribution [3], leading to performance degradation. To mitigate these challenges, existing approaches require access to source training data and labels for complete re-training [4, 5], continual learning [3] or test-time training [6], which is costly and time-consuming.
In this work, we address these limitations by introducing a novel TTA-based method, namely Think Twice before Adaptation (T2A), which enhances pre-trained DF detectors without requiring access to source training data or labels. Our approach achieves two key objectives: (1) enhanced resilience through dynamic adaptation to unknown postprocessing techniques; and (2) improved generalization to new samples from unknown distributions. While current TTA approaches commonly employ Entropy Minimization (EM) as the adaptation objective, solely relying on EM can result in confirmation bias caused by overconfident predictions [7] and model collapse [8]. To this end, in T2A, we design a novel Uncertainty-aware Negative Learning adaptation objective with noisy pseudo-labels, allowing the model to explore alternative options (i.e., other classes in the classification problem) rather than becoming overly confident in potentially incorrect predictions. For better adaptation, we incorporate Focal Loss [9] into the negative learning (NL) objective to dynamically prioritize crucial samples and propose a gradients masking technique that updates crucial model parameters whose gradients align with those of BatchNorm layers.
@@ 2.1 Deepfake Detection
DF detection approaches are often formulated as a binary classification problem that automatically learns discriminative features from large-scale datasets [10]. Existing approaches can be classified into three categories based on their inputs: (i) Spatial-based approaches that operate directly on pixel-level features [4, 11, 12], (ii) Frequency-based approaches that analyze generation artifacts in the frequency domain [13, 14], and (iii) Hybrid approaches that integrate both pixel and frequency domain information within a unified method [15, 16]. Recent advances have improved the cross-dataset generalization of DF detectors by employing data augmentation (DA) strategies [4, 17], synthesis techniques [5, 5], continual learning [3], meta-learning and one-shot test-time training [6].
Compared to existing methods, our T2A offers advantages: (1) T2A enables DF detectors to be adapted to test data without access to source data (e.g., OST [6] requires source data for adaptation); (2) T2A does not rely on any DA or synthesis techniques to extend the diversity of data; (3) Not only enhance the generalization, T2A also improves the resilience of DF detectors to unknown postprocessing techniques. Additionally, our method is orthogonal to these works [18, 19, 20], which require pre-training on joint datasets (physical and digital attacks) and do not adapt during inference.
@@ 2.2 Test-time Adaptation (TTA)
TTA approaches only require access to the pre-trained model from the source domain for adaptation [21]. Unlike source-free domain adaptation approaches [22], which require access to the entire target dataset, TTA enables online adaptation to the arrived test samples.
TENT [23] and MEMO [24] optimized batch normalization (BN) statistics from the test batch through EM. LAME [25] adapted only the model’s output probabilities by minimizing Kullback–Leibler divergence between the model’s predictions and optimal nearby points’ vectors. Several methods have studied TTA in continuously changing environments. CoTTA [26] implemented weight and augmentation averaging to mitigate error accumulation, while EATA [27] developed an efficient entropy-based sample selection strategy for model updates. Inspired by parameter-efficient fine-tuning, VIDA [28] used high-rank adapters to handle domain shifts. However, these methods solely rely on EM as the learning principle, which can present two issues: (1) Confirmation bias: EM greedily pushes for confident predictions on all samples, even when predictions are incorrect [7], leading to overconfident yet incorrect predictions; and (2) Model Collapse: EM tends to cause model collapse, where the model predicts all samples to the same class, regardless of their true labels [8]. The model collapse phenomenon is particularly problematic in DF detection, where the inherent bias toward dominant fake samples in training data [29] can exacerbate the collapse.
@@ 2.3 Negative Learning
Supervised learning or positive learning (PL) directly maps inputs to their corresponding labels. However, when labels are noisy, PL can lead models to learn incorrect patterns. Negative learning (NL) [30] addresses this challenge by training networks to identify which classes an input does not belong to. Several loss functions have been proposed by leveraging this concept: NLNL [30] combines sequential PL and NL phases, while JNPL [31] proposes a single-phase approach through joint optimization of enhanced NL and PL loss functions. Recent work has further integrated NL principles with normalization techniques [32] to transform active losses into passive ones [33].
@@ 3 Generation Artifacts Analysis
Artifacts in DFs generated by Generative Adversarial Examples (GANs), which emerge from the upsampling operations in the GANs pipeline, can be revealed in the frequency domain through Discrete Fourier Transform (DFT) [14]. In this section, we demonstrate postprocessing techniques can completely obscure these artifacts presented in DF samples, leading to performance degradation of DF detectors.
@@
This property (Proof in Appendix A) is particularly important for understanding why the upsampling operation leaves artifacts in the frequency domain [34]. For an image $x(\cdot,\cdot)$ convolved with a kernel $c(\cdot,\cdot)$, the output $y(\cdot,\cdot)$ in the spatial domain and its frequency domain form can be expressed as:
@@
$\tilde{X}(u,v)=\begin{cases}X(u,v),&\;u\in[0,M-1],v\in[0,N-1]\\ X(u-M,v),&\;u\in[M,2M-1],v\in[0,N-1]\\ X(u,v-N),&\;u\in[0,M-1],v\in[N,2N-1]\\ X(u-N,v-N),&\;u\in[M,2M-1],v\in[N,2N-1]\end{cases}$ (6)
@@
However, these spectral artifacts exhibit vulnerability to various postprocessing operations [2]. As shown in Figure 1(b), the GAN-generated image displays distinctive checkerboard artifacts in its frequency spectrum, but they undergo substantial modifications when subjected to different postprocessing operations (Figures 1(c)-(d)). The magnitude of these artifacts’ obscurity correlates directly with the intensity of the applied postprocessing operations, as demonstrated in Figure 3 (Appendix B). Furthermore, the empirical analysis presented in Figure 2 of Appendix B shows that the performance of existing DF detectors tends to drop significantly when encountering unseen postprocessing techniques with increasing intensities.
@@ 4.1 Problem Definition
Given a DF detector $f:\mathcal{X}\rightarrow\mathbb{R}^{2}$ parameterized by $\theta$ is well-trained on the training data $\mathcal{D}^{train}=\{(x_{i},y_{i})\}^{N^{train}}_{i=1}\sim P^{train}(x,y)$, where $x\in\mathcal{X}$ is the input and $y\in\mathcal{Y}=\{0,1\}$ is the target label, our goal is to online update parameters $\theta$ of $f$ on mini-batches $\{\mathcal{B}_{1},\mathcal{B}_{2},\dots\}$ of the test stream $\mathcal{D}^{test}=\{(x_{j},y_{j})\}^{N^{test}}_{j=1}\sim P^{test}(x,y)$. Note that, in the online TTA setting, $P^{train}(x,y)$ and $\{y_{j}\}$ are unavailable, and the knowledge learned in previously seen mini-batches could be accumulated for adaptation to the current mini-batch [21]. In this work, we consider online TTA in two challenging scenarios of DF detection:
@@ 4.3.1 Uncertainty Modelling with Noisy Pseudo-Labels
where $\tau\in[0,1]$ denotes the classification threshold. Rather than implicitly trusting the model’s initial predictions, we enable the model to "doubt" its predictions by introducing noisy pseudo-labels.
@@ 4.3.2 Noise-tolerant Negative Loss Function
From Positive to Negative Learning. Negative learning (NL) enables the model to be taught with a lesson that """"this input image does not belong to this complementary label" [30]. In our work, converting from pseudo-labels to noisy versions is equivalent to transforming from positive to negative learning, facilitating the DF model to re-think that "this input image might not belong to this real (fake)/fake (real) label"""".
Noise-tolerant Negative Loss Function. Inspired by existing works [35, 32, 36], we start from the fact that any loss function can be robust to noisy labels through a simple normalization operation:
@@
As shown in Figure 5 (Appendix C.4.4), given a normalized loss function with pseudo label $\mathcal{L}_{norm}(x,\hat{y})$, our normalized negative loss function $\mathcal{L}_{nn}(x,\tilde{y})$ with noisy pseudo-label is the opposite of $\mathcal{L}_{norm}(x,\hat{y})$.
Prior research by [32, 33] has indicated that the normalized loss function suffers from the underfitting problem. This problem is particularly critical in the TTA context where the model only "sees" a few samples during inference. To address this challenge, we incorporate the passive loss function $\mathcal{L}_{p}$ [33] into TTA, leading to our NTNL which can effectively help the model to adapt to noisy pseudo-labels:
@@ 4.4 Uncertain Sample Prioritization
To identify which samples should be prioritized during adaptation, we propose a dynamic prioritization strategy that focuses on uncertain samples (i.e., low confidence). Our intuition here is that lower-confidence samples require the model to be considered more carefully. Specifically, we incorporate Focal Loss [9] into the NTNL function (Eq. 12). Formally, the loss function $\ell(x,\tilde{y})$ is now defined:
@@ 4.5 Gradients Masking
BatchNorm (BN) adaptation [37] is widely used in existing TTA approaches [27, 23]. BN is a crucial layer that normalizes each feature $z$ during training: $y=\varrho*\left(\frac{(z-\mu^{b})}{\sigma^{b}}\right)+\vartheta$, where $\mu^{b}$ and $\sigma^{b}$ are batch statistics, and $\varrho$, $\vartheta$ are learnable parameters. After training, $\mu^{ema}$ and $\sigma^{ema}$, which are estimated over the whole training dataset via exponential moving average (EMA) [37], are used during inference. When $P^{train}(x,y)\neq P^{test}(x,y)$, BN adaptation replaces EMA statistics ($\mu^{ema}$, $\sigma^{ema}$) with statistics computed from test mini-batches ($\hat{\mu}^{b}$, $\hat{\sigma}^{b}$). However, this approach is limited by only updating BN layer parameters.
@@
where $\psi$ is a threshold to control the selection of parameters for updating. This technique brings more capacity for adaptation as more model parameters are updated compared to approaches that only update BN parameters during inference [27, 23].
@@ 5.1.1 Datasets and modeling
We use Xception [38] as the source model, which as commonly used as the backbone in DF detectors. The training set is FaceForensics++ (FF++) [39]. To evaluate the adaptability of our T2A method, we use six more datasets at inference time, including CelebDF-v1 [40], CelebDF-v2 [40], DeepFakeDetection (DFD) [41], DeepFake Detection Challenge Preview (DFDCP) [42], UADFV [43], and FaceShifter (FSh) [44]. The dataset implementations are provided by [45] and more details are described in Appendix C.

### v2

@@ Abstract
Deepfake (DF) detectors face significant challenges when deployed in real-world environments, particularly when encountering test samples deviated from training data through either postprocessing manipulations or distribution shifts. We demonstrate postprocessing techniques can completely obscure generation artifacts presented in DF samples, leading to performance degradation of DF detectors. To address these challenges, we propose Think Twice before Adaptation (T2A), a novel online test-time adaptation method that enhances the adaptability of detectors during inference without requiring access to source training data or labels. Our key idea is to enable the model to explore alternative options through an Uncertainty-aware Negative Learning objective rather than solely relying on its initial predictions as commonly seen in entropy minimization (EM)-based approaches. We also introduce an Uncertain Sample Prioritization strategy and Gradients Masking technique to improve the adaptation by focusing on important samples and model parameters. Our theoretical analysis demonstrates that the proposed negative learning objective exhibits complementary behavior to EM, facilitating better adaptation capability. Empirically, our method achieves state-of-the-art results compared to existing test-time adaptation (TTA) approaches and significantly enhances the resilience and generalization of DF detectors during inference.
@@ 1 Introduction
Recently, Generative Artificial Intelligence (GenAI) has been used to generate DFs for malicious purposes, such as impersonation [^1] Finance worker pays out \$currency-dollar\\$\$25 million after video call with deepfake ‘chief financial officer’ and disinformation spread [^2] AI-faked images of Donald Trump’s imagined arrest swirl on Twitter, raising concerns about privacy and security. Several DF detection approaches have been proposed to mitigate these negative impacts Nguyen-Le et al. (2024a). Despite advances, deploying these systems in real-world environments presents two critical challenges. First, in practice, adversaries can strategically apply previously unknown postprocessing techniques to DF samples at inference time, completely obscuring the generation artifacts Corvi et al. (2023) and successfully bypassing detection systems. Second, real-world applications are frequently exposed to test samples drawn from distributions that deviate substantially from the training data distribution Pan et al. (2023), leading to performance degradation. To mitigate these challenges, existing approaches require access to source training data and labels for complete re-training Ni et al. (2022); Shiohara and Yamasaki (2022), continual learning Pan et al. (2023) or test-time training Chen et al. (2022), which is costly and time-consuming.
In this work, we address these limitations by introducing a novel TTA-based method, namely Think Twice before Adaptation (T2A), which enhances pre-trained DF detectors without requiring access to source training data or labels. Our approach achieves two key objectives: (1) enhanced resilience through dynamic adaptation to unknown postprocessing techniques; and (2) improved generalization to new samples from unknown distributions. While current TTA approaches commonly employ Entropy Minimization (EM) as the adaptation objective, solely relying on EM can result in confirmation bias caused by overconfident predictions Zhang et al. (2024) and model collapse Niu et al. (2023). To this end, in T2A, we design a novel Uncertainty-aware Negative Learning adaptation objective with noisy pseudo-labels, allowing the model to explore alternative options (i.e., other classes in the classification problem) rather than becoming overly confident in potentially incorrect predictions. For better adaptation, we incorporate Focal Loss Ross and Dollár (2017) into the negative learning (NL) objective to dynamically prioritize crucial samples and propose a gradients masking technique that updates crucial model parameters whose gradients align with those of BatchNorm layers.
@@ 2.1 Deepfake Detection
DF detection approaches are often formulated as a binary classification problem that automatically learns discriminative features from large-scale datasets Nguyen-Le et al. (2024b). Existing approaches can be classified into three categories based on their inputs: (i) Spatial-based approaches that operate directly on pixel-level features Ni et al. (2022); Cao et al. (2022), (ii) Frequency-based approaches that analyze generation artifacts in the frequency domain Liu et al. (2021); Frank et al. (2020), and (iii) Hybrid approaches that integrate both pixel and frequency domain information within a unified method Liu et al. (2023b). Recent advances have improved the cross-dataset generalization of DF detectors by employing data augmentation (DA) strategies Ni et al. (2022); Yan et al. (2024), synthesis techniques Shiohara and Yamasaki (2022), continual learning Pan et al. (2023), meta-learning and one-shot test-time training Chen et al. (2022).
Compared to existing methods, our T2A offers advantages: (1) T2A enables DF detectors to be adapted to test data without access to source data (e.g., OST Chen et al. (2022) requires source data for adaptation); (2) T2A does not rely on any DA or synthesis techniques to extend the diversity of data; (3) Not only enhance the generalization, T2A also improves the resilience of DF detectors to unknown postprocessing techniques. Additionally, our method is orthogonal to these works Fang et al. (2024); Liu et al. (2024); He et al. (2024), which require pre-training on joint datasets (physical and digital attacks) and do not adapt during inference.
@@ 2.2 Test-time Adaptation (TTA)
TTA approaches only require access to the pre-trained model from the source domain for adaptation Liang et al. (2024). Unlike source-free domain adaptation approaches Li et al. (2024), which require access to the entire target dataset, TTA enables online adaptation to the arrived test samples.
TENT Wang et al. (2020) and MEMO Zhang et al. (2022) optimized batch normalization (BN) statistics from the test batch through EM. LAME Boudiaf et al. (2022) adapted only the model’s output probabilities by minimizing Kullback–Leibler divergence between the model’s predictions and optimal nearby points’ vectors. Several methods have studied TTA in continuously changing environments. CoTTA Wang et al. (2022) implemented weight and augmentation averaging to mitigate error accumulation, while EATA Niu et al. (2022) developed an efficient entropy-based sample selection strategy for model updates. Inspired by parameter-efficient fine-tuning, VIDA Liu et al. (2023a) used high-rank adapters to handle domain shifts. However, these methods solely rely on EM as the learning principle, which can present two issues: (1) Confirmation bias: EM greedily pushes for confident predictions on all samples, even when predictions are incorrect Zhang et al. (2024), leading to overconfident yet incorrect predictions; and (2) Model Collapse: EM tends to cause model collapse, where the model predicts all samples to the same class, regardless of their true labels Niu et al. (2023). The model collapse phenomenon is particularly problematic in DF detection, where the inherent bias toward dominant fake samples in training data Layton et al. (2024) can exacerbate the collapse.
@@ 2.3 Negative Learning
Supervised learning or positive learning (PL) directly maps inputs to their corresponding labels. However, when labels are noisy, PL can lead models to learn incorrect patterns. Negative learning (NL) Kim et al. (2019) addresses this challenge by training networks to identify which classes an input does not belong to. Several loss functions have been proposed by leveraging this concept: NLNL Kim et al. (2019) combines sequential PL and NL phases, while JNPL Kim et al. (2021) proposes a single-phase approach through joint optimization of enhanced NL and PL loss functions. Recent work has further integrated NL principles with normalization techniques Ma et al. (2020) to transform active losses into passive ones Ye et al. (2023).
@@ 3 Generation Artifacts Analysis
Artifacts in DFs generated by Generative Adversarial Examples (GANs), which emerge from the upsampling operations in the GANs pipeline, can be revealed in the frequency domain through Discrete Fourier Transform (DFT) Frank et al. (2020). In this section, we demonstrate postprocessing techniques can completely obscure these artifacts presented in DF samples, leading to performance degradation of DF detectors.
@@
This property (Proof in Appendix A) is particularly important for understanding why the upsampling operation leaves artifacts in the frequency domain Ojha et al. (2023). For an image $x(\cdot,\cdot)$ convolved with a kernel $c(\cdot,\cdot)$, the output $y(\cdot,\cdot)$ in the spatial domain and its frequency domain form can be expressed as:
@@
$\tilde{X}(u,v)=\begin{cases}X(u,v),&\;u\in[0,M-1],v\in[0,N-1]\\ X(u-M,v),&\;u\in[M,2M-1],v\in[0,N-1]\\ X(u,v-N),&\;u\in[0,M-1],v\in[N,2N-1]\\ X(u-N,v-N),&\;u\in[M,2M-1],v\in[N,2N-1]\end{cases}$
@@
However, these spectral artifacts exhibit vulnerability to various postprocessing operations Corvi et al. (2023). As shown in Figure 1(b), the GAN-generated image displays distinctive checkerboard artifacts in its frequency spectrum, but they undergo substantial modifications when subjected to different postprocessing operations (Figures 1(c)-(d)). The magnitude of these artifacts’ obscurity correlates directly with the intensity of the applied postprocessing operations, as demonstrated in Figure 3 (Appendix B). Furthermore, the empirical analysis presented in Figure 2 of Appendix B shows that the performance of existing DF detectors tends to drop significantly when encountering unseen postprocessing techniques with increasing intensities.
@@ 4.1 Problem Definition
Given a DF detector $f:\mathcal{X}\rightarrow\mathbb{R}^{2}$ parameterized by $\theta$ is well-trained on the training data $\mathcal{D}^{train}=\{(x_{i},y_{i})\}^{N^{train}}_{i=1}\sim P^{train}(x,y)$, where $x\in\mathcal{X}$ is the input and $y\in\mathcal{Y}=\{0,1\}$ is the target label, our goal is to online update parameters $\theta$ of $f$ on mini-batches $\{\mathcal{B}_{1},\mathcal{B}_{2},\dots\}$ of the test stream $\mathcal{D}^{test}=\{(x_{j},y_{j})\}^{N^{test}}_{j=1}\sim P^{test}(x,y)$. Note that, in the online TTA setting, $P^{train}(x,y)$ and $\{y_{j}\}$ are unavailable, and the knowledge learned in previously seen mini-batches could be accumulated for adaptation to the current mini-batch Liang et al. (2024). In this work, we consider online TTA in two challenging scenarios of DF detection:
@@ 4.3.1 Uncertainty Modelling with Noisy Pseudo-Labels
where $\tau\in[0,1]$ denotes the classification threshold. Rather than implicitly trusting the model’s initial predictions, we enable the model to ”doubt” its predictions by introducing noisy pseudo-labels.
@@ 4.3.2 Noise-tolerant Negative Loss Function
From Positive to Negative Learning. Negative learning (NL) enables the model to be taught with a lesson that """"this input image does not belong to this complementary label” Kim et al. (2019). In our work, converting from pseudo-labels to noisy versions is equivalent to transforming from positive to negative learning, facilitating the DF model to re-think that ”this input image might not belong to this real (fake)/fake (real) label"""".
Noise-tolerant Negative Loss Function. Inspired by existing works Zhou et al. (2021); Ma et al. (2020); Ghosh et al. (2017), we start from the fact that any loss function can be robust to noisy labels through a simple normalization operation:
@@
As shown in Figure 4 (Appendix C.2), given a normalized loss function with pseudo label $\mathcal{L}_{norm}(x,\hat{y})$, our normalized negative loss function $\mathcal{L}_{nn}(x,\tilde{y})$ with noisy pseudo-label is the opposite of $\mathcal{L}_{norm}(x,\hat{y})$.
Prior research by Ma et al. (2020); Ye et al. (2023) has indicated that the normalized loss function suffers from the underfitting problem. This problem is particularly critical in the TTA context where the model only ”sees” a few samples during inference. To address this challenge, we incorporate the passive loss function $\mathcal{L}_{p}$ Ye et al. (2023) into TTA, leading to our NTNL which can effectively help the model to adapt to noisy pseudo-labels:
@@ 4.4 Uncertain Sample Prioritization
To identify which samples should be prioritized during adaptation, we propose a dynamic prioritization strategy that focuses on uncertain samples (i.e., low confidence). Our intuition here is that lower-confidence samples require the model to be considered more carefully. Specifically, we incorporate Focal Loss Ross and Dollár (2017) into the NTNL function (Eq. 12). Formally, the loss function $\ell(x,\tilde{y})$ is now defined:
@@ 4.5 Gradients Masking
BatchNorm (BN) adaptation Schneider et al. (2020) is widely used in existing TTA approaches Niu et al. (2022); Wang et al. (2020). BN is a crucial layer that normalizes each feature $z$ during training: $y=\varrho*\left(\frac{(z-\mu^{b})}{\sigma^{b}}\right)+\vartheta$, where $\mu^{b}$ and $\sigma^{b}$ are batch statistics, and $\varrho$, $\vartheta$ are learnable parameters. After training, $\mu^{ema}$ and $\sigma^{ema}$, which are estimated over the whole training dataset via exponential moving average (EMA) Schneider et al. (2020), are used during inference. When $P^{train}(x,y)\neq P^{test}(x,y)$, BN adaptation replaces EMA statistics ($\mu^{ema}$, $\sigma^{ema}$) with statistics computed from test mini-batches ($\hat{\mu}^{b}$, $\hat{\sigma}^{b}$). However, this approach is limited by only updating BN layer parameters.
@@
where $\psi$ is a threshold to control the selection of parameters for updating. This technique brings more capacity for adaptation as more model parameters are updated compared to approaches that only update BN parameters during inference Niu et al. (2022); Wang et al. (2020).
@@ 5.1.1 Datasets and modeling
We use Xception Chollet (2017) as the source model, which as commonly used as the backbone in DF detectors. The training set is FaceForensics++ (FF++) Rossler et al. (2019). To evaluate the adaptability of our T2A method, we use six more datasets at inference time, including CelebDF-v1 Li et al. (2020b), CelebDF-v2 Li et al. (2020b), DeepFakeDetection (DFD) Google (2019), DeepFake Detection Challenge Preview (DFDCP) Dolhansky (2019), UADFV Li et al. (2018), and FaceShifter (FSh) Li et al. (2020a). The dataset implementations are provided by Yan et al. (2023) and more details are described in Appendix C.